# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v12

@@ Abstract
Coal extraction was an influential economic activity in interwar Japan. In the initial stage, coal mines used not only males but also females as the core workforce in the pits. However, the innovation of labor-saving technologies and the renewal of traditional extraction methodology induced institutional change through the revision of labor regulations on female miners in the early 1930s. This dramatically changed the mines as the place where skilled males were the principal miners engaged in the underground works. I investigate the impact of coal mining on regional growth and assess how the institutional change induced by the labor regulations affected its process. By linking the location information of mines with registration- and census-based statistics, I found that coal mines led to remarkable population growth. The labor regulations did not stagnate but accelerated the local population growth as they forced female miners to exit from the labor market and form their families. The regulations prohibited risky underground works by female miners. This reduction in occupational hazards also improved early-life mortality via the mortality selection mechanism in utero.
@@ 1 Introduction
How institution influences the economy has attracted wider attention (Hall and Jones 1999; Acemoglu et al. 2001). Institutional interventions in the labor market may impede development as they could prevent the labor market and structural adjustments (Freeman 1993; Fernández-Villaverde 2018). Cross-country evidence indicates that stronger regulation on labor markets could be associated with lower labor force participation and higher unemployment (Botero et al. 2004). [^1] Djankov et al. (2002) using the dataset on the regulation of entry by start-up firms in 85 countries provide evidence that more robust regulation of entry of start-up firms could be associated with higher corruption, larger unofficial economies, and worse quality of public and private goods. A different view by Nickell and Laynard (1999) argues that labor institutions such as labor unions are more likely to be associated with economic performance than institutional regulation in European countries. Evidence based on the domestic variations in the labor regulation shows that pro-worker labor regulation decreases the economic activities in the formal manufacturing sector in India (Besley and Burgess 2004). However, a specific type of regulation, such as unemployment insurance, may improve risk-sharing and reduce the transaction cost in job search (Acemoglu and Shimer 1999). Therefore, understanding what labor regulations could hinder (or improve) economic performance is still an influential research agenda.
@@
In this paper, I investigate how coal mining impacts regional economic development in interwar Japan and how the institutional change concerning female labor patterns induced by the technology shocks influenced it. In prewar Japan, the coal mining sector relied not only on male miners but also on female miners. However, the innovation of labor-saving technologies such as coal cutters and conveyors and the renewal of traditional extraction methodology induced institutional change through the revision of labor regulations on female miners in the interwar period. It dramatically changed the mining place from where female and male manual miners worked in the pits to where skilled male miners were the principal workers. Such an institutional change could have impacted the regional economy as it increased the gender wage gap.
@@
I found that the coal mine leads to population growth, indicating the regional development in the mining area. The revised labor regulations did not stagnate the local population growth but accelerated it because the institutional change due to the regulations had forced female miners to exit from the labor market and form their families. Since the revised regulation prohibited risky work in the pits by the female miners, it substantially reduced their occupational hazards. This led to the declines in overall female mortality and improvements in the early-life mortality rates via the mortality selection mechanism in utero.
@@
This paper adds to the literature on the role of the state and institutions during economic development from the perspective of female labor regulations (Botero et al. 2004; Besley and Burgess 2004). Importantly, it is the first study investigating the impacts of institutional change due to the regulations of female workers in the resource extraction sector. Coal is necessary for industrialization because broad economic activities heavily rely on coal consumption. I show that technology-induced labor regulations reduced female labor force participation and accelerated the local population growth through the family formation channel. This institutional change mitigated the occupational hazards for female miners and improved the early-life health status in the local economy. In this light, my findings shed new light on the role of gender-specific labor regulations in understanding the mechanisms behind regional economic growth.
@@
This study is relevant to the broader literature on the economic impacts of resource extractions. Predictions regarding agglomeration theory imply that resource abundance can increase the demand for local labor in the mining sector and related industries, thereby causing an influx of population and agglomeration (Rosenthal and Strange 2004; Duranton and Puga 2004; Lederman and Maloney 2007). [^2] Induced agglomeration can improve efficiency in the local labor markets through, for instance, better matching and learning, causing spillover effects in the non-mining sectors (Black et al. 2005; Bjørnland and Thorsrud 2016). See Corden and Neary (1982); Eastwood and Venables (1982); Corden (1984); van Wijnbergen (1984); Neary and van Wijnbergen (1984) for earlier theoretical studies on the impact of resource discoveries. However, an increase in the resource extraction industry may stagnate the development of non-resource extraction industries by reallocating resources from the non-resource tradable sector (Sachs and Warner 1999, 2001). [^3] The wealth effect mechanism suggests that resource abundance increases the demand for commodities, thereby reallocating resources from the tradable to non-tradable sector and increasing the imports of tradable. This can reduce labor supply, leading to greater earnings in the local economy. The direct effect on the relative demand of extraction firms and workers is higher prices of personal and business services. Other political-economic factors that may disrupt the growth of the non-resource sector are rent-seeking and corruption. Caselli and Guy (2009, 2013) summarize these theoretical implications. See Papyrakis and Gerlagh (2007) for empirical evidence using US states’ data. In fact, there are still debates on whether proximity to coal mines is necessary for regional economic growth during Europe’s industrialization. For example, a recent study by Fernihough and O’Rourke (2020) found that after the mid-eighteenth century, European cities closer to coalfields had a larger population than other cities. [^4] Similarly, a few economic history studies have argued for the importance of better access to coalfields in subsequent industrial development (Mathias 2001; Pollard 1981). By contrast, Mokyr (1977) observed that the distribution of coal supplies did not contribute to industrialization. My main results support the former assertion by showing that the coal mines led to population growth in the local economy. Importantly, this study adds to this literature to analyze the mechanism behind population growth due to resource extraction. [^5] Although studies have analyzed the impact of resource abundance on population growth, the mechanisms underlying this relationship have been understudied. See for example Black et al. (2005); Michaels (2010); Fernihough and O’Rourke (2020). I show that while the population growth in the mining area initially occurred due to internal migration, the miner households formed families in response to the institutional change.
@@
This paper also expands our understanding of the relationship between mining and regional human capital accumulation. A pivotal risk of mining is pollution, which can increase the health costs in resource-abundant locations and decrease agricultural productivity (Aragón and Rud 2015; von der Goltz and Barnwal 2019). To the best of my knowledge, however, the potential health risks of mining during historical industrialization have yet to be evaluated. [^6] The economic history studies have predominantly focused on the health costs of air pollution. For example, Heblich et al. (2021) and Beach and Hanlon (2017) found that extensive coal use during historical economic growth in Britain increased infant mortality rates and decreased the average adult height. Relatedly, Hanlon (2020) found that coal-related pollution disturbed the growth of cities in the late nineteenth century in the UK. Another strand of the literature investigates the adverse impact of using lead pipes on mortality rates (Grönqvist et al. 2020). I find that while there is little evidence on the air and water pollution pathways, occupational hazards for female miners lead to mortality selection in utero, which increases early-life mortality. Regulations for female miners are suggested to mitigate the mortality sorting mechanisms before births. My finding sheds light on the importance of occupational hazards in evaluating the economic activities of the resource extraction industry, which has been neglected in the literature. In addition, it provides suggestive evidence that labor regulations may improve the human capital of their offspring by improving early-life health conditions (e.g., Almond et al. 2018).
@@
Finally, this study utilizes data constructed using census-based statistics. [^7] Most studies have focused on large-scale, particularly gold, mines and utilized the location information combined with survey datasets (Aragón and Rud 2013, 2015; Kotsadam and Tolonen 2016). This allows me to eliminate the risks caused by selection bias in the statistical inference. This study also exploits specific geological strata given in nature as an instrumental variable, which offers robust estimates for evaluating the impacts of the coal mines on the regional economy. Among the literature, this study provides the first evidence of a past-developing country in Asia, given that research has mainly focused on European countries, the US, Latin America, and African countries. [^8] See, for example, Domenech (2008); Fernihough and O’Rourke (2020) for European countries, Black et al. (2005); Michaels (2010) for the US, Sachs and Warner (1999); Aragón and Rud (2013) for Latin America, and Aragón and Rud (2015); Benshaul-Tolonen (2019) for African countries.
@@
The remainder of the study is organized as follows. Section 2 briefly overviews the historical context. Section 3 introduces the conceptual framework for this paper. Section 4 describes the data and Section 5 summarizes empirical analysis. Section 6 concludes.
@@ 2.1 Coal Mining Industry
The Mining Code of 1892 allowed private companies to have mining concessions, accelerating the installation of modern technologies in mines and the development of Japan’s mining sector. For instance, winding engines were installed in the coal and copper mines in the 1890s, and the flame-smelting process was adopted in gold and silver mines in the 1900s (Ishii 1991, pp. 220–221). Before the First World War (WWI), the coal industry had developed as an export labor-intensive industry. However, WWI increased the average wage rates of miners, reducing the comparative advantage of the Japanese coal industry. Figure 1 illustrates the development of the coal industry and shows the negative shock in the early 1920s. This motivated coal mining firms to reduce the number of miners and adopt new extraction machines to improve firm productivity. [^9] See Sumiya (1968, pp. 387–393) for details on the adoption process of these machines. Mining firms reallocate resources to relatively productive mines to improve their average productivity. Consequently, average labor productivity in the coal mining sector increased from the 1920s, particularly in the 1930s (Okazaki 2021). [^10] Generally, during the interwar period, the mine and manufacturing industries experienced steady capital investments, including investments in production facilities and machines, and public investments (Nakamura 1971, pp. 138–143). Consequently, the relative contribution of manufacturing and mining, which provided energy and resources to manufacturing, was the largest at 42-51% among all sectors during this period (Minami 1994, p.90). Figure 1 shows that total coal production increased during the interwar period.
@@
In the early 1930s, coal accounted for approximately $55$% of total mineral output in Japan (Mining Bureau, Ministry of Commerce and Industry 1934, p.44). The large share of coal production indicates that coal extraction was an influential economic activity at that time. The higher wage rates of miners have motivated peasants to move into the extraction industry, shifting the local industrial structure from the primary to the production sector. For instance, the average daily wage of female miners working in the pits was one and a half to two times higher than those of the female workers in the silk and spinning industries (Nishinarita 1985, pp. 84–85). Besides the original inhabitants, therefore, mining workers would include a certain proportion of migrants. This may have influenced the structure of both the marriage market and fertility in the local economy surrounding mines. However, quantitative evidence of these changes in the regional economy is still lacking in literature. [^11] Most research on mines in prewar Japan comprises business history studies that investigated the features of the management and industrial organization in coal mining companies (Ishii 1991, pp. 220-233). Tanaka (1984) provided a comprehensive review of the works in the Japanese coal mining industry.
@@ 2.2 Mining Technology, Occupational Hazards, and Regulations of Female Miners
By the end of the 19th century, a large part of mines had finished mechanizing the processes involved in hauling, such as drainage pumps and hoisting machines. However, this did not mean that the miner’s workload related to hauling operations was reduced. Instead, such mechanization in the main shafts had increased the workload of miners extracting and transporting coal around the face (Sumiya 1968, pp. 310–312). Moreover, the environment in the pit was generally poor. The air was polluted with dust, oxygen levels were low, and there was a risk of carbon monoxide poisoning. There was also a high fire risk from dynamite blasting and spontaneous combustion of gas and coal. Accidents caused by falling rocks occurred frequently. Available statistics indicate that the average number of accidents per year in the pits between $1917$ and $1926$, causing injuries and deaths of miners, was $142,595$, of which $61,112$ were from falling rocks. Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by falling rocks, but they still occurred $311$ times per year on average between $1917$ and $1926$. [^12] These statistics are from Japan Mining Association (1928, p. table.6).
@@
Notably, while miners included single male workers, many of them were married couples as miners usually worked in teams of two to three (Sumiya 1968, pp. 300–308). The male skilled miner, called saki yama, extracted coal at the face; meanwhile, the female miner, named ato yama, brought coal to the coal wagon at the gangway through steep pits using bamboo baskets called sura (Sumiya 1968, p. 319). As coal is a bulky resource, the heavy work burden on pregnant women could have increased the risk of adverse pregnancy outcomes (Ahmed and Jaakkola 2007). Miners worked in pits with dirty air containing particulates, which eventually damaged their lungs and increased the risk of respiratory diseases such as tuberculosis (Sumiya 1968, pp.  410–411). [^13] A health survey for miners found that the coal miners had the highest prevalence among all types of miners at that time, which suggests how hard the work in the coal mines was (Japan Mining Association 1928, p. 3). Relatedly, a growing body of medical literature has revealed that particulates and radionuclides from coal increase the risk of stillbirths and premature deaths due to respiratory diseases in mothers (Landrigan et al. 2017; Lin et al. 2013).
@@
An important event that changed the harsh labor pattern of female miners was an institutional change that occurred in the early 1930s. Following the International Labor Conference after WWI, the Social Affairs Bureau held a consultation meeting with miners in the 1920s to revise the Miners’ Labor Assistance Regulations of 1916. Initially, mining companies, who relied heavily on female miners for production, opposed the Social Affairs Bureau’s insistence on prohibiting women from working underground and late-night. However, the renewal of mining methods driven by new extraction technologies changed the attitudes of the mining companies in the late 1920s. In fact, the transition from the traditional room-and-pillar mining method to the longwall method was accompanied by the adoption of technologies such as coal cutters and conveyors. Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of those machines (Nishinarita 1985, p. 91). Specifically, the conveyors could bring extracted coals from pits to gangway further efficiently than the female miners did, adding strong incentive to the mining companies to prohibit the underground works by the female miners (Tanaka and Ogino 1977, p. 79). [^14] Online Appendix A.1 provides finer details in the process of this mechanization. Online Appendix A.2 provides cross-sectional evidence on the replacement of female miners with the use of machines.At last, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work by female miners. Since the grace period for enforcement of the regulations lasted until 1933, the number of female miners gradually declined from around 1930. This change induced by mechanization was commonly observed in Japanese coal mines (Tanaka 1984, p. 392). [^15] A detailed chronology of technology adoptions in each coal mine in Chikuhō coalfield confirms this movement (Compilation Committee of the Chronology of Chikuhō Coal Mining Industry 1973).
@@
Panel A of Table 1 shows the evidence that the number of female miners in the pits dropped from 1930 to 1933 when the revised regulation was enacted in September. Despite this, the full enforcement of the revised labor regulations did not eliminate female miners altogether because females were still allowed to work in the pits of the thin layer and were primarily responsible for coal selection (sentan) and other out-of-pit labor. In the late 1930s, more than $15,000$ women still worked in the mines, representing approximately $10$% of all the miners measured. In this light, female miners played an important role in the coal mines throughout the interwar period. The enforcing regulations, however, impacted the market values of female workers in the mining area. Column 1 of Panel B in Table 1 presents evidence that the relative wage rate of female miners dropped after the enforcement. In fact, while there was a decreasing trend in the relative wage rate under the mechanization process in the 1920s, the difference between 1930 and 1933 shows the largest decline in the relative wage rate ($0.66$ v. $0.48$). This reduction is not a macroeconomic trend around the mining region, given that the relative wage rate in the agricultural sector is stable during the same period (Column 2 of Panel B). The rate for in-pit workers dropped from $0.78$ to $0.62$, whereas that for out-pit workers was relatively stable from $0.49$ to $0.45$. The mean difference before and after the enforcement year is statistically significant only in the in-pit workers (bottom of Panel B). This suggests that regulating in-pit female miners did not improve the relative wage rate of the out-pit female miners. Therefore, although prohibiting female workers in the risky deep layers should mitigate the occupational hazards for female miners, it shall lower the relative wages for females in the mining area.
@@
Notes: The statistics on the miners in this table are from the Labor Statistics Field Survey Reports that surveyed mines employing 50 or more miners. % share of female miners in Panel A shows the percentage share of females to the total number of miners. The relative wage rate in Panel B is the female average daily wage divided by the male average daily wage. The overall ratios for the miners between 1927 and 1936 (column 1) are recalculated using the number of miners worked in and outside the pits and their daily wages. The mean difference in the wage ratio around 1933 and $t$-statistic $p$-value for the mean difference are listed at the bottom of Panel B.
Sources: Statistics are from Statistics Bureau of the Cabinet (1926, pp. 5; 32), Statistics Bureau of the Cabinet (1930, pp. 6–7; 26; 30), Statistics Bureau of the Cabinet (1932b, pp. 6–7; 26; 30; 38), Statistics Bureau of the Cabinet (1936, pp. 41; 64; 68; 76), and Statistics Bureau of the Cabinet (1937, pp. 45; 78; 79). The average daily wages for the daily and annual agricultural laborers are obtained from Umemura et al. (1966, p. 107). The description of the survey subject is from Statistics Bureau of the Cabinet (1932a, p. 1) and Statistics Bureau of the Cabinet (1937, p. 1).

### v14

@@ Abstract
Coal extraction was an influential economic activity in interwar Japan. Initially, coal mines employed both males and females as the workforce in the pits. However, the innovation of labor-saving technologies and the renewal of traditional extraction methodology induced institutional change. This was manifested by the revision of labor regulations affecting female miners in the early 1930s. This dramatically changed the mining workplace, making skilled males the principal miners engaged in underground work. This paper investigates the impact of coal mining on regional growth and assesses how the institutional changes induced by the amended labor regulations affected its processes. By linking the mines’ location information with both registration and census-based statistics, it was found that coal mines led to remarkable population growth. The labor regulations accelerated local population growth by forcing female miners to exit the labor market and form families. The regulations prohibited female miners from risky underground work. This reduction in occupational hazard also improved early-life mortality via the mortality selection mechanism in utero.
@@ 1 Introduction
How institutions influence the economy has attracted wide attention (Hall and Jones 1999; Acemoglu et al. 2001). Institutional interventions in the labor market may impede development by preventing labor market and structural adjustments (Freeman 1993; Fernández-Villaverde 2018). International evidence indicates that strong regulation over labor markets can be associated with lower labor force participation and higher unemployment (Botero et al. 2004). [^1] Djankov et al. (2002) used a dataset on start-up firms’ entry regulations from 85 countries. They found that more robust regulation of start-up firms’ entry can be associated with higher corruption, larger unofficial economies, and worse quality of public and private goods. A different view by Nickell and Laynard (1999) argues that, in European countries, labor institutions such as labor unions are more likely to be associated with economic performance than institutional regulation. Evidence based on the domestic variations in labor regulation shows that in India, pro-worker labor regulation decreases economic activity in the formal manufacturing sector (Besley and Burgess 2004). However, a specific type of regulation, such as unemployment insurance, may improve risk-sharing and reduce the transaction cost of job searching (Acemoglu and Shimer 1999). Therefore, understanding the type of labor regulations that could hinder (or improve) economic performance is still an influential research agenda.
@@
This paper investigates how coal mining impacted regional development during interwar Japan, and how it was influenced by the institutional change concerning female labor patterns induced by technology shocks. In prewar Japan, the coal mining sector relied on both male and female miners. However, during the interwar period there was innovation through labor-saving technologies such as coal cutters and conveyors combined with the renewal of traditional extraction methodology. Simultaneously there was revision of labor regulations affecting female miners. It dramatically changed the mining workplace from one where both female and male manual miners worked in the pits to one where skilled male miners became the principal workers. This institutional change impacted the regional economy by increasing the gender wage gap.
@@
I found that the presence of coal mines leads to population growth, indicating regional development in the mining area. The revised labor regulations did not stagnate the local population growth but accelerated it. This was because the institutional change due to the regulations forced female miners to exit the labor market and establish families. Since the revised regulation prohibited female miners from risky work in the pits, it substantially reduced their occupational hazards. This led to declines in overall female mortality and improvements in the early-life mortality rates via the mortality selection mechanism in utero.
@@
This paper contributes to the literature on the role of the state and institutions during economic development from the perspective of female labor regulations (Botero et al. 2004; Besley and Burgess 2004). Importantly, it is the first study investigating the impact of institutional change due to changes in the regulations governing female workers in the resource extraction sector. Coal is necessary for industrialization because wide variety of economic activities heavily rely on coal consumption. I showed that technology-induced labor regulations reduced female labor force participation and accelerated the local population growth through the family formation channel. This institutional change mitigated the occupational hazards for female miners and improved the early-life health status in the local economy. From this perspective, the findings shed new light on the role of gender-specific labor regulations in understanding the mechanisms behind regional development.
@@
This study is relevant to the broader literature on the economic impact of resource extraction. Predictions regarding agglomeration theory imply that resource abundance can increase the demand for local labor in the mining sector and related industries, thereby causing an influx of population and agglomeration (Rosenthal and Strange 2004; Duranton and Puga 2004; Lederman and Maloney 2007). [^2] Induced agglomeration can improve efficiency in the local labor markets through, for instance, better matching and learning. This can have spillover effects in the non-mining sectors (Black et al. 2005; Bjørnland and Thorsrud 2016). See Corden and Neary (1982); Eastwood and Venables (1982); Corden (1984); van Wijnbergen (1984); Neary and van Wijnbergen (1984) for earlier theoretical studies on the impact of resource discoveries. However, growth of the resource extraction industry may stagnate the development of non-resource extraction industries by reallocating resources from the non-resource tradable sector (Sachs and Warner 1999, 2001). [^3] The wealth effect mechanism suggests that resource abundance increases the demand for commodities, thereby reallocating resources from the tradable to non-tradable sector and increasing the imports of tradables. This can reduce labor supply, leading to greater earnings in the local economy. The direct effect on the relative demand of extraction firms and workers is higher prices of personal and business services. Other political-economic factors that may disrupt the growth of the non-resource sector are rent-seeking and corruption. Caselli and Guy (2009, 2013) summarize these theoretical implications. See Papyrakis and Gerlagh (2007) for empirical evidence using United States data. There is still debate on whether proximity to coal mines was necessary for regional development during Europe’s industrialization. A recent study by Fernihough and O’Rourke (2020) found that after the mid-eighteenth century, European cities closer to coalfields had larger populations than other cities. [^4] Similarly, some economic history studies have argued for the importance of better access to coalfields in subsequent industrial development (Mathias 2001; Pollard 1981). By contrast, Mokyr (1977) observed that the distribution of coal supplies did not contribute to industrialization. My main results support the former assertion by showing that the coal mines led to population growth in the local economy. Importantly, this study contributes to the literature analyzing the mechanism behind population growth due to resource extraction. [^5] Although studies have analyzed the impact of resource abundance on population growth, the mechanisms underlying this relationship have been understudied. See for example Black et al. (2005); Michaels (2010); Fernihough and O’Rourke (2020). I demonstrate that while the population growth in the mining area initially occurred due to internal migration, the miner households formed families in response to the institutional change.
@@
This paper also expands understanding of the relationship between mining and regional human capital accumulation. A pivotal risk of mining is pollution, which can increase the health costs in resource-abundant locations and decrease agricultural productivity (Aragón and Rud 2015; von der Goltz and Barnwal 2019). Academic consensus is that the potential health risks of mining during historical industrialization are still to be evaluated. [^6] Economic history studies have focused on the health costs of air pollution. Heblich et al. (2021) and Beach and Hanlon (2017) found that extensive coal use during historical economic growth in Britain increased infant mortality rates and decreased the average adult’s height. Relatedly, Hanlon (2020) found that coal-related pollution disturbed the growth of United Kingdom (UK)cities in the late nineteenth century. Another strand of the literature investigates the adverse impact of using lead pipes on mortality rates (Grönqvist et al. 2020). I found that while there is little evidence on the air and water pollution pathways, occupational hazards for female miners led to mortality selection in utero, which increased early-life mortality. Regulations for female miners are suggested to mitigate the mortality sorting mechanisms before births. The finding sheds light on the importance of occupational hazards in evaluating the economic activities of the resource extraction industry, which has been neglected in the literature. In addition, it provides evidence suggesting that labor regulations may improve the human capital of their offspring by improving early-life health conditions (e.g., Almond et al. 2018).
@@
Finally, this study utilizes data constructed using census-based statistics. [^7] Most studies have focused on large-scale mines, particularly gold, and utilized location information combined with survey datasets (Aragón and Rud 2013, 2015; Kotsadam and Tolonen 2016). This enables the elimination of risks caused by selection bias in the statistical inference. This study also exploits specific geological strata given in nature as an instrumental variable. This offers robust estimates for evaluating the impact of coal mines on the regional economy. In the literature, this study provides the first evidence from a developing Asian country in the past. This significant given that earlier research has mainly focused on European countries, the US, Latin America, and African countries. [^8] See, for example, Domenech (2008); Fernihough and O’Rourke (2020) for European countries, Black et al. (2005); Michaels (2010) for the US, Sachs and Warner (1999); Aragón and Rud (2013) for Latin America, and Aragón and Rud (2015); Benshaul-Tolonen (2019) for African countries.
@@
The remainder of the study is organized as follows. Section 2 briefly overviews the historical context. Section 3 introduces the conceptual framework for this paper. Section 4 describes the data and Section 5 summarizes empirical analysis. Section 6 concludes the study.
@@ 2.1 Coal Mining Industry
The Mining Code of 1892 allowed private companies to have mining concessions, accelerating the installation of modern technologies in mines and the development of Japan’s mining sector. Winding engines were installed in coal and copper mines in the 1890s, and the flame-smelting process was adopted in gold and silver mines in the 1900s (Ishii 1991, pp. 220–221). Before the First World War (WWI), however, the Japanese coal industry had developed as a labor-intensive export industry. Then, the outbreak of WWI increased miners’ average wage rates, reducing the coal industry’s comparative advantage. Figure 1 illustrates the development of the coal industry and shows the negative shock in the early 1920s. This motivated coal mining firms to reduce the number of miners and adopt new extraction machines to improve productivity. [^9] See Sumiya (1968, pp. 387–393) for details on the machinery adoption process. Mining firms reallocated resources to relatively productive mines to improve their average productivity. Consequently, average labor productivity in the coal mining sector increased from the 1920s, particularly in the 1930s (Okazaki 2021). [^10] Generally, during the interwar period, the mining and manufacturing industries experienced steady capital investment, particularly in production facilities and machines, and public investments (Nakamura 1971, pp. 138–143). Consequently, the relative contribution of manufacturing and mining, which provided energy and resources to manufacturing, was the largest at 42-51% among all economic sectors during this period (Minami 1994, p.90). Figure 1 shows that total coal production increased during the interwar period. [^11] As the moderate decline in the early 1930s confirmed, the impact of the depression was not substantial compared to the other countries. See Online Appendix A.1 for details.
@@
In the early 1930s, coal accounted for approximately $55$% of total mineral output in Japan, meaning that coal extraction was an influential economic activity at that time (Mining Bureau, Ministry of Commerce and Industry 1934, p.44). The higher wage rates of miners motivated peasants to move into the extraction industry, shifting the local industrial structure from the agricultural to the production sector. An example, the average daily wage of female miners working in the pits was one and a half to two times higher than those of the female workers in the silk and spinning industries (Nishinarita 1985, pp. 84–85). Besides the original inhabitants, therefore, mining workers would include a certain proportion of migrants. This may have influenced the structure of both the marriage market and fertility in the local economy surrounding mines. However, quantitative evidence of these changes in the regional economy is still lacking in literature. [^12] Most research on mines in prewar Japan comprises business history studies that investigated the features of the management and industrial organization in coal mining companies (Ishii 1991, pp. 220-233). Tanaka (1984) provided a comprehensive review of the operations of the Japanese coal mining industry.
@@ 2.2 Mining Technology, Occupational Hazards, and Regulations of Female Miners
By the end of the 19th century, a large proportion of mines had completed mechanizing the hauling processes. This was done by installing equipment such as drainage pumps and hoisting machines. However, this did not mean that the miners’ workload related to hauling operations was reduced. Instead, such mechanization in the main shafts increased the workload of miners extracting and transporting coal around the coalface (Sumiya 1968, pp. 310–312). Moreover, the environment in the pit was generally poor. The air was polluted by dust, oxygen levels were low, and there was a risk of carbon monoxide poisoning. There was also a high fire risk from dynamite blasting and spontaneous combustion of gas and coal. Accidents caused by rockfall occurred frequently. Available statistics indicate that the average number of accidents per year in the pits between $1917$ and $1926$, causing miners’ injuries and deaths, was $142,595$, of which $61,112$ were from rockfall. Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by rockfall, but they still occurred $311$ times per year on average between $1917$ and $1926$. [^13] These statistics are from Japan Mining Association (1928, p. table.6).
@@
Notably, while miners included single male workers, many of them were married couples. This was because miners usually worked in teams of two to three (Sumiya 1968, pp. 300–308). The male skilled miner, called saki yama, extracted coal at the face; meanwhile, the female miner, named ato yama, carried that coal to the coal wagon at the gangway. She traveled through steep pits carrying bamboo baskets called sura (Sumiya 1968, p. 319). As coal is a bulky resource, the heavy work burden on pregnant women could have increased the risk of adverse pregnancy outcomes (Ahmed and Jaakkola 2007). Miners worked in pits with dirty air containing particulates, which eventually damaged their lungs and increased the risk of respiratory diseases such as tuberculosis (Sumiya 1968, pp.  410–411). [^14] A miners’ health survey found that coal miners had the highest respiratory disease prevalence among all types of miners at that time, which indicated how hard the work in the coal mines was (Japan Mining Association 1928, p. 3). Relatedly, a growing body of medical literature has revealed that particulates and radionuclides from coal increase the risk of stillbirths and premature deaths due to mothers’ respiratory diseases (Landrigan et al. 2017; Lin et al. 2013).
@@
An important event that changed the harsh labor pattern experienced by female miners was an institutional change occurring in the early 1930s. Following the International Labor Conference after WWI, the Social Affairs Bureau held a consultation with mining companies in the 1920s to revise the Miners’ Labor Assistance Regulations of 1916. Initially, mining companies, who relied heavily on female miners for production, opposed the Social Affairs Bureau’s insistence on prohibiting women from working underground and late-night. However, the renewal of mining methods driven by new extraction technologies changed the attitudes of the mining companies in the late 1920s. The transition from the traditional room-and-pillar mining method to the longwall method was accompanied by the adoption of technologies such as coal cutters and conveyors. Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of those machines (Nishinarita 1985, p. 91). Specifically, the conveyors brought extracted coals from pits to the gangway more efficiently than the female miners, strongly incentivizing the mining companies to prohibit underground work by female miners (Tanaka and Ogino 1977, p. 79). [^15] Online Appendix A.2 provides finer details in this mechanization process. Online Appendix A.3 provides cross-sectional evidence on the replacement of female miners by machines. Finally, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work by female miners. As the grace period for enforcement of the regulations lasted until 1933, the number of female miners gradually declined from around 1930. This change induced by mechanization was commonly observed in Japanese coal mines (Tanaka 1984, p. 392). [^16] A detailed chronology of technology adoptions in each coal mine in Chikuhō coalfield confirms this movement (Compilation Committee of the Chronology of Chikuhō Coal Mining Industry 1973).
@@
Panel A of Table 1 demonstrates the evidence that the number of female miners in the pits reduced from 1930 to 1933 when the revised regulation was enacted. The full enforcement of the revised labor regulations did not eliminate female miners altogether because females were still allowed to work in the thin layer pits. They were primarily responsible for coal selection (sentan) and other out-of-pit labor. In the late 1930s, more than $15,000$ women still worked in the mines, representing approximately $10$% of all the miners measured. From this perspective, female miners played an important role in the coal mines throughout the interwar period. The enforcement of regulations, however, impacted the market values of female workers’ skills in the mining area. Column 1 of Panel B in Table 1 presents evidence that female miners’ relative wage rate dropped after the enforcement. While there was a decreasing trend in the relative wage rate during the mechanization process in the 1920s, the difference between 1930 and 1933 shows the largest decline ($0.66$ v. $0.48$). This reduction was not a macroeconomic trend around the mining region, given that the relative wage rate in the agricultural sector was stable during the same period (Column 2 of Panel B). Moreover, the rate for in-pit workers reduced from $0.78$ to $0.62$, whereas that for out-pit workers was relatively stable from $0.49$ to $0.45$. The mean difference before and after the enforcement year is statistically significant only in the in-pit workers’ category (bottom of Panel B). This indicates that regulating in-pit female miners did not improve the relative wage rate of the out-pit female miners. Therefore, although prohibiting female workers in the risky deep layers should mitigate the occupational hazards for female miners, it shall lower the relative wages for females in the mining area.
@@
Notes: The statistics on the miners in this table are drawn from the Labor Statistics Field Survey Reports that surveyed mines employing 50 or more miners. % share of female miners in Panel A shows the percentage share of females relative to the total number of miners. The relative wage rate in Panel B is the female average daily wage divided by the male average daily wage. The overall ratios for the miners between 1927 and 1936 (column 1) are recalculated using the number of miners working inside and outside of the pits and their daily wages. The mean difference in the wage ratio around 1933 and $t$-statistic $p$-values for the mean difference are listed at the bottom of Panel B.
Sources: Statistics are from Statistics Bureau of the Cabinet (1926, pp. 5; 32), Statistics Bureau of the Cabinet (1930, pp. 6–7; 26; 30), Statistics Bureau of the Cabinet (1932b, pp. 6–7; 26; 30; 38), Statistics Bureau of the Cabinet (1936, pp. 41; 64; 68; 76), and Statistics Bureau of the Cabinet (1937, pp. 45; 78; 79). The average daily wages for the daily and annual agricultural laborers are obtained from Umemura et al. (1966, p. 107). The survey subject description is from Statistics Bureau of the Cabinet (1932a, p. 1) and Statistics Bureau of the Cabinet (1937, p. 1).